OpenStudy (anonymous):

-tan^2 (x) + sec^2 (x) = 1

terenzreignz (terenzreignz):

Are we supposed to prove this?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

-tan^2x+1+tan^2x =1

terenzreignz (terenzreignz):

Well then, I suggest you start with the well known Pythagorean identity.... \[\Large \sin^2(x) + \cos^2(x) = 1\]

OpenStudy (anonymous):

sec^2x can be written as 1+tan^2x and t

OpenStudy (anonymous):

both the negative na positiv tab^2 x gets cancelled

hartnn (hartnn):

or draw a triangle and use pythagoras theorem.

OpenStudy (anonymous):

@madrockz so how would it look like , I'm more of a visual learner

hartnn (hartnn):

lets not use pythagorean identity to prove a pythagorean identity...

hartnn (hartnn):

draw a triangle, right angles label its sides and angles

hartnn (hartnn):

right angled triangle*

OpenStudy (anonymous):

Oh I see now ,thanks everyone